OpenStudy (yanni):

|6u+3| =15 solve for u

OpenStudy (anonymous):

The rule of thumb when solving absolute value is divide into 2 cases: 6u + 3 = 15 6u + 3 = -15 Can you continue?

OpenStudy (yanni):

6u=-15-3 6u=-18 u=-3?

OpenStudy (alexwee123):

there are two solutions for "u" look at Chlorophyll's answer again

OpenStudy (yanni):

which are correct

OpenStudy (yanni):

which is correct*

OpenStudy (alexwee123):

there are two correct answers

OpenStudy (yanni):

kk thx

OpenStudy (anonymous):

Are you fine? Do you me to check your results?

OpenStudy (anonymous):

Keep in mind, absolute value has 2 solutions! Here u = 2 and u = -3